Hello,
This homework question is from a worksheet, not a book:
F=2E+10
E 1 2 3 4 7 10 100
F
I need to fill in the numbers for line F.
My answer is 12 14 16 18 24 30 210
My answer assumes that 2E is 2 multiplied by each number on line E. Is this correct? Thanks in advance.

Answer by checkley71(8403)

BECAUSE F=2E+10 THEN WHEN E=1 F=2(1)+10 OR F=2+10 OPR F=12
WHEN E=2 THEN F=2(2)+10 OR F=4+10 OR F=14
WHEN E=100 THEN F=2(100)+10 OR F=200+10 OR F=210
YOU'RE ON THE RIGHT TRACK